# Exploring the Potential of Encoder-free Architectures in 3D LMMs

### v1

@@ 2.2 LLM-embedded Semantic Encoding
Knowledge Distillation Loss. We select the powerful Uni3D-L (Zhou et al., 2023) as the teacher encoder, input the point cloud into the 3D encoder, and obtain the output feature $F_{\text{teacher}}\in\mathbb{R}^{M\times D_{2}}$. The Mean Squared Error (MSE) between the LLM output tokens $F_{\text{student}}$ and $F_{\text{teacher}}$ is computed to align $F_{\text{student}}$ as closely as possible to $F_{\text{teacher}}$, thereby transferring the knowledge embedded in the 3D encoder to the LLM. By obtaining additional supervision from the Uni3D, the LLM better captures the complex structures in the point cloud data, as displayed in Figure 3 (d). The objective function can be defined as
@@
  • The point cloud self-supervised learning losses generally benefit the encoder-free 3D LMM. Compared to previous experimental results, where the GPT scores for the classification and captioning tasks are 49.11% and 45.39%, the self-supervised losses bring about the significant improvements. This is because the self-supervised learning loss forces transformations on the complex point clouds through certain task design. This encourages the LLM to not simply memorize specific point cloud data but to learn the underlying geometric relationships and high-level semantic information.
  • Among the self-supervised learning losses, the Masked Modeling Loss demonstrates the strongest performance improvement. It achieves GPT-4 scores of 49.5% and 47.35% for classification and captioning tasks, respectively. The application of the masked modeling to the point features facilitates the embedding of high-level semantics from point clouds into the LLM. However, the mask ratio is directly proportional to the training optimization difficulty, and increasing it from 30% to 60% results in a performance degradation. In addition, explicitly reconstructing point patches is not as effective as masked modeling in capturing the critical features of the input, but it does help the LLM learn the complex patterns within point clouds. Knowledge Distillation Loss falls short compared to the first two losses. Finally, Contrastive Loss, which fails to extract the detailed semantics, achieves the lowest performance.
@@
Enel-7B 50.92 48.61 49.31 3.88 7.20 12.50 55.00 42.70
Hybrid Semantic Loss. Based on the experimental results above, we propose the self-supervised learning loss specifically designed for the encoder-free 3D LMM—Hybrid Semantic Loss, as showcased in Figure 3 (e). We apply a masking ratio $r$ to randomly mask point tokens from the token embedding. The masked tokens and the corresponding patches are referred to as $\{F_{\text{mask}_{i}}\}_{i=1}^{M*r}$ and $\{G_{\text{mask}_{i}}\}_{i=1}^{M*r}$, respectively. The remaining tokens are denoted as$\{F_{\text{vis}_{i}}\}_{i=1}^{M*(1-r)}$ and $\{G_{\text{vis}_{i}}\}_{i=1}^{M*(1-r)}$. For the masked portion, we adopt masked modeling, and for the visible portion, we use the reconstruction strategy. The inverse modeling process is described in Appendix C.1. Considering the autoregressive nature of the LLM and the unordered attribute of point clouds, we directly concatenate learnable tokens $\{F_{\text{learn}_{i}}\}_{i=1}^{M*r}$ to the end of $F_{\text{vis}}$, replacing the masked tokens. After passing point tokens through the LLM, we compute the MSE between $F_{\text{learn}}$ and $F_{\text{mask}}$. The visible features $F_{\text{vis}}$ are transformed into $G_{\text{pred}}$, and the $L_{2}$ Chamfer distance is computed between $G_{\text{pred}}$ and $G_{\text{vis}}$. These two are added to the original cross-entropy loss with coefficients all equal to 1. This approach not only embeds high-level semantics into the LLM but also ensures geometric consistency throughout the point cloud learning process. With a mask ratio of 30%, it achieves 52.00% and 48.51% for the classification and captioning tasks. Adding positional encodings to the learnable and visible tokens at each layer of the LLM further yields improvements of +1% and +0.34%, respectively, enhancing the perception of spatial positional information.
@@ 2.3 Hierarchical Geometry Aggregation
3D encoders are designed with specific structures tailored for point clouds, such as local-to-global hierarchy (Zhang et al., 2022a) for exploring the geometric structure of the point cloud. However, in encoder-free architectures, the LLM itself does not have an explicit local modeling module. The self-attention mechanism is intended for modeling global interactions. Therefore, building upon the proposed Hybrid Semantic Loss, we explore in the instruction tuning stage how to enable the LLM to actively perceive 3D local details and complement the learned global semantics. To this end, we propose the Hierarchical Geometry Aggregation strategy in the LLM early layers.
@@
Then we perform $l-1$ iterations of geometry aggregation, resulting in $\{F_{\text{agg}_{i}}^{l}\}_{i=1}^{M/2^{l}}$. To ensure that the LLM fully extracts the local information, we choose to perform further semantic modeling using $H$ LLM layers after aggregation operations. This allows the model to learn the interactions between local information while preventing the loss of fine-grained geometric details. Subsequently, we perform $l$ iterations of geometry propagation. Following the approach of PointNet++ (Qi et al., 2017), we propagate the aggregated features $F_{\text{agg}}^{l}$ from the local center points to their surrounding $k$ neighboring points, generating $\{F_{\text{pro}_{i}}^{1}\}_{i=1}^{M/2^{(l-1)}}$. After $l$ iterations, we obtain point tokens of length $M$, which are then processed by the remaining $L-2l-(H+1)$ layers.
@@
  • The best performance is achieved when $l$ is set to 1 and the performance decreases as $l$ increases. As observed in Table 4, performing single aggregation and propagation operation achieves 48.86% and 52.5% performance on the captioning and classification tasks, respectively, while additional operations lead to a performance drop of 3%-4%. This is because as $l$ increases, repeated aggregation operations progressively simplify spatial relationships, causing the loss of fine-grained geometric structures in the point clouds.
  • Compared to setting $H$ to 4 or 8, the highest performance is achieved when $H$ is set to 2. It reaches 53.5% and 49.13% on the classification and captioning tasks, respectively. The excessive number of LLM layers between aggregation and propagation can lead to the oversmoothing of the aggregated local information, resulting in the loss of local structural details.
  • The introduction of the gated self-attention mechanism effectively enhances performance, achieving the GPT score of 55% in the classification task and 50.92% in the captioning task. The adaptive control of attention output ensures that global contextual information is utilized only when necessary, preventing it from disrupting local geometric structures. Additionally, it allows the model to adjust to different tasks.
@@ 3 Results and Visualization
Results. In Table 5, on the Objaverse benchmark (Deitke et al., 2023), Enel-7B achieves the GPT score of 50.92% on the 3D object captioning task, setting a new state-of-the-art (SOTA) performance. In traditional metrics, Sentence-BERT and SimCSE reaches 48.61% and 49.31%, respectively, comparable to the ShapeLLM-13B. For the 3D object classification task, Enel-7B outperformes prior encoder-based 3D LMMs, achieving a GPT score of 55%. Given the same training dataset as PointLLM, these results validate the effectiveness of our proposed LLM-embedded Semantic Encoding and Hierarchical Geometry Aggregation strategies for the encoder-free architecture. Additionally, on the 3D-VQA task of the 3D MM-Vet dataset (Qi et al., 2024), despite the lack of spatial and embodied interaction-related data in the training set, Enel achieves the GPT score of 42.7%, surpassing PointLLM-7B by 1.5%.
@@ 4 Conclusion
Appendix A Related Work
Encoder-free Vision-Language Models. Traditional vision-language models (VLMs) often rely on vision encoders to extract visual features before processing them with language models, integrating image encoders like CLIP (Radford et al., 2021) and DINO V2 (Oquab et al., 2023). However, recent efforts have explored encoder-free VLMs for their simplicity. Approaches like (ChameleonTeam, 2024; Xie et al., 2024) use VQ tokenizers (Esser et al., 2021) or linear projection layers (Diao et al., 2024a; Chen et al., 2024a) to represent images. Fuyu-8B (Bavishi et al., 2023), a pure decoder-only model, directly processes image patches through linear projections, handling high-resolution images but showing only average performance. The EVE (Diao et al., 2024b) eliminates the need for a separate vision encoder by bridging vision-language representation within a unified decoder and enhancing visual recognition capabilities through additional supervision.
Figure 6: Variants of Point Cloud Self-Supervised Learning Losses. (a) The Variant of Masked Modeling Loss, (b) The Variant of Reconstruction Loss, (c) The Variant of Hybrid Semantic Loss.
Appendix B Experimental Settings
Implementation Details. We use the LLaMA model (Touvron et al., 2023) as our LLM backbone, with the 7B Vicuna-v1.1 (Chiang et al., 2023) checkpoint as the default setting. In the token embedding layer, the point cloud is first processed by a linear layer to expand its dimension from 6 to 288. The input point cloud initially consists of 8192 points, followed by three iterations of farthest point sampling (FPS), reducing the size to 512, 256, and 128, respectively. After each FPS operation, k-Nearest Neighbors (k-NN) is applied with a cluster size of 81. And geometric features are extracted using triangular encoding, followed by linear layers that progressively increase the dimension to 576, 1152, and 2304. Finally, the projection layer maps the features to the LLM dimension of 4096. In the pre-training stage, we unfreeze the first four LLM layers. Within the LLM-embedded Semantic Encoding strategy, Hybrid Semantic Loss applies masked modeling to 30% of the tokens and reconstructs the patches for the remaining 70% visible tokens. In the instruction tuning stage, we apply geometric aggregation in the second LLM layer, reducing the number of point tokens from 128 to 64. MaxMean pooling is used to retain more information. After passing through two LLM layers, the geometric aggregation is applied in the fifth layer to restore the point size count to 128.
@@ Appendix B Experimental Settings
Table 6: Ablation Experiments. We begin the ablation experiments by changing the single configuration of the module from Enel. $\Psi$ represents a mask ratio of 60%, while $\Phi$ represents a mask ratio of 30%. For Hybrid Semantic Loss, the subscript $patch$ and $feat$ represent the masked modeling target, while the reconstruction target is the corresponding $feat$ and $patch$. $l$ represents the number of aggregation and propagation operations. $H$ refers to the LLM layers between $l$ aggregation and $l$ propagation operations. $O$ refers to the LLM layer between two individual aggregation or propagation operations.
Model Cap Cls
@@
Enel-7B 50.92 48.61 49.31 3.88 7.20 12.50 55.00
– Hybrid Semantic Loss 47.19 48.07 48.31 3.46 7.41 11.84 50.61
Hybrid Semantic LosspatchΦ 49.05 48.82 49.20 4.01 7.25 12.38 52.20
Hybrid Semantic LosspatchΨ 48.96 48.38 49.00 3.66 6.97 11.98 52.00
Hybrid Semantic LossfeatΨ 49.63 48.00 48.62 3.78 6.88 12.33 51.50
– gate mechanism 49.26 48.41 48.93 3.71 7.12 12.47 53.50
l=2,H=2,O=0 48.81 48.10 48.57 3.70 6.99 12.01 51.50
l=2,H=4,O=0 49.02 48.47 48.61 3.65 7.10 12.31 52.00
l=2,H=2,O=2 48.96 47.96 48.89 3.80 7.05 12.55 52.00
l=2,H=4,O=2 49.58 48.70 48.84 3.84 7.56 12.76 53.00
Appendix C More Experiments
C.1 Variants of Point Cloud Self-Supervised Learning Losses.
@@ Appendix C More Experiments
C.2 More Ablation Experiments
@@
We begin the ablation experiments starting from the Enel, which is the reverse order compared to the experiments in the main text, as showcased in Tabel 6
@@
The Effects of LLM-embedded Semantic Encoding Strategy. In the Table 6, on the basis of Enel, removing the Hybrid Semantic Loss during the pre-training stage significantly degrades performance. The GPT-4 score for the captioning task drops from 50.92% to 47.19%, and the GPT-4 score for the classification task decreases to 50.61%. This is because the proposed self-supervised learning function for point clouds effectively captures the detailed structures and high-level semantics of the point clouds.
@@
Based on Enel, we find that setting the mask ratio in the Hybrid Semantic Loss to 30% consistently yields better results than 60%. Additionally, the configuration where the masked token part predicts features while the visible token part reconstructs patches outperforms the reverse setting—where the masked token part predicts patches and the visible token part reconstructs features. This phenomenon can be explained as follows: a mask ratio of 30% retains critical information while facilitating the model to effectively utilize the visible tokens to derive the masked parts. When the mask ratio is set too high, the model fails to learn the global context knowledge adequately. Moreover, when the masked token part is tasked with predicting features, the model focuses on learning the high-level context semantics, while the patch reconstruction aids in accurately capturing low-level details. In contrast, when the masked token part predicts patches, the model becomes excessively dependent on local features during the process of semantic reconstruction.
@@
The Effects of Hierarchical Geometry Aggregation Strategy. After removing the gating mechanism in the self-attention of the aggregation operation, the performance drops to 49.26% and 53.50% on the captioning and classification tasks, respectively. The gating mechanism helps the model to adaptively filter information, allowing it to focus on more discriminative features. Without the dynamic adjustment to focus on different parts of the input, the generated text from the LLM lacks accuracy and coherence, leading to a decrease in performance.
@@
The performance generally degrades with an increasing number of aggregation and propagation operations. This degradation can be attributed to the progressive loss of local geometric details through repeated aggregation operations, while multiple propagation operations typically rely on interpolation which tends to amplify high-frequency noise. We observe that increasing the number of LLM layers between the final aggregation operation and the first propagation operation leads to improved performance. This suggests that cascaded aggregation operations necessitate deeper architectural capacity for high-level feature abstraction, as insufficient network depth may lead to degradation of hierarchical representations. Furthermore, the presence of LLM layers between each aggregation or propagation operation enhances performance by allowing the model to process and transform compressed information. Through self-attention mechanisms, these intermediate layers can recapture and restore details lost during the aggregation process.
Appendix D Model Output

### v2

@@ 3.2 LLM-embedded Semantic Encoding
Knowledge Distillation Loss. We select the powerful Uni3D-L (Zhou et al., 2023) as the teacher encoder, input the point cloud into the 3D encoder, and obtain the output feature $F_{\text{teacher}}\in\mathbb{R}^{M\times D_{2}}$. The Mean Squared Error (MSE) between the LLM output tokens $F_{\text{student}}$ and $F_{\text{teacher}}$ is computed to align $F_{\text{student}}$ as closely as possible to $F_{\text{teacher}}$, thereby transferring the knowledge embedded in the 3D encoder to the LLM. By obtaining additional supervision from the Uni3D, the LLM better captures the complex structures in the point cloud data, as displayed in Figure 3 (d). The objective function is defined as
@@
  • The point cloud self-supervised losses generally benefit the encoder-free 3D LMM. Compared to previous experimental results, where the GPT scores for the classification and captioning tasks are 49.11% and 45.39%, the self-supervised losses bring about the significant improvements. This is because the self-supervised learning loss forces transformations on the complex point clouds through certain task design. This encourages the LLM to not simply memorize specific point cloud data but to learn the underlying geometric relationships and high-level semantic information.
  • Among the self-supervised learning losses, the Masked Modeling Loss demonstrates the strongest performance improvement. It achieves GPT-4 scores of 49.5% and 47.35% for classification and captioning tasks, respectively. The application of the masked modeling to the point features facilitates the embedding of high-level semantics from point clouds into the LLM. However, a higher mask ratio increases training difficulty, with 60% performing worse than 30%. In addition, explicitly reconstructing point patches helps capture complex structures and critical details in point clouds. Knowledge Distillation Loss falls short compared to the first two losses. Finally, Contrastive Loss, which fails to extract the detailed semantics, achieves the lowest performance.
Hybrid Semantic Loss. Based on the experimental results above, we propose the self-supervised learning loss specifically designed for the encoder-free 3D LMM—Hybrid Semantic Loss, as showcased in Figure 3 (e). We apply a masking ratio $r$ to randomly mask point tokens from the token embedding. The masked tokens and the corresponding patches are referred to as $\{F_{\text{mask}_{i}}\}_{i=1}^{M*r}$ and $\{G_{\text{mask}_{i}}\}_{i=1}^{M*r}$, respectively. The remaining tokens are denoted as$\{F_{\text{vis}_{i}}\}_{i=1}^{M*(1-r)}$ and $\{G_{\text{vis}_{i}}\}_{i=1}^{M*(1-r)}$. Considering the autoregressive nature of the LLM and the unordered attribute of point clouds, we directly concatenate learnable tokens $\{F_{\text{learn}_{i}}\}_{i=1}^{M*r}$ to the end of $F_{\text{vis}}$, replacing the masked tokens. For the masked portion, we adopt masked modeling, and for the visible portion, we use the reconstruction strategy. After passing point tokens through the LLM, we compute the MSE between $F_{\text{learn}}$ and $F_{\text{mask}}$. The visible features $F_{\text{vis}}$ are transformed into $G_{\text{pred}}$, and the $L_{2}$ Chamfer distance is computed between $G_{\text{pred}}$ and $G_{\text{vis}}$. These two are added to the original cross-entropy loss with coefficients all equal to 1. This approach not only embeds high-level semantics into the LLM but also ensures geometric consistency throughout the point cloud learning process. With a 30% mask ratio and per-layer positional encoding of point tokens, it achieves 53.00% and 48.85% on the classification and captioning tasks, respectively. The inverse modeling process is described in Appendix B.1.
@@
Enel-7B 50.98 48.65 49.36 3.85 7.20 12.60 55.10 43.10
Enel-13B 52.87 48.85 50.03 3.72 7.25 12.31 55.30 48.20
@@ 3.3 Hierarchical Geometry Aggregation
3D encoders are designed with specific structures tailored for point clouds, such as local-to-global hierarchy (Zhang et al., 2022) for exploring the geometric structure of the point cloud. However, in encoder-free architectures, the LLM itself does not have an explicit local modeling module. The self-attention mechanism is intended for modeling global interactions. Therefore, building upon the proposed Hybrid Semantic Loss, we explore in the instruction tuning stage how to enable the LLM to actively perceive 3D local details and complement the learned global semantics. To this end, we propose the Hierarchical Geometry Aggregation strategy.
@@
Then we perform $l-1$ iterations of geometry aggregation, resulting in $\{F_{\text{agg}_{i}}^{l}\}_{i=1}^{M/2^{l}}$. To ensure that the LLM fully extracts the local information, we choose to perform further semantic modeling using $H$ LLM layers after aggregation operations. This allows the model to learn the interactions between local information while preventing the loss of fine-grained geometric details. Subsequently, from the $L$th layer, we perform $l$ iterations of geometry propagation. Following the approach of PointNet++ (Qi et al., 2017), we propagate the aggregated features $F_{\text{agg}}^{l}$ from the local center points to their surrounding $k$ neighboring points, generating $\{F_{\text{pro}_{i}}^{1}\}_{i=1}^{M/2^{(l-1)}}$. After $l$ iterations, we obtain point tokens of length $M$, which are then processed by the remaining LLM layers.
@@
  • The best performance is achieved when $l$ is set to 3. As observed in Table 4, performing triple aggregation and propagation operations achieves 48.92% and 52.60% performance on the captioning and classification tasks, respectively. Fewer aggregation layers $l$ limit local geometric capture, while too many overly simplify spatial relationships. Setting $l=3$ achieves a balanced modeling of local and global structures.
  • Compared to setting $H$ to 4 or 8, the highest performance is achieved when $H$ is set to 2. It reaches 53.58% and 49.20% on the classification and captioning tasks, respectively. The excessive number of LLM layers between aggregation and propagation can lead to the oversmoothing of the aggregated local information, resulting in the loss of local structural details.
  • The gated self-attention mechanism effectively improves performance, reaching 55.10% and 50.98% on classification and captioning tasks, respectively. The adaptive control of attention output ensures that global contextual information is utilized only when necessary, preventing it from disrupting local geometric structures. Additionally, it allows the model to adjust to different tasks.
Figure 6: Enel Response Examples. We demonstrate that Enel is capable of providing accurate responses across different types of tasks, such as captioning and question answering, by effectively addressing a wide range of objects, including race cars, buildings, insects, and others.
@@ 4 Results and Visualization
Results. In Table 5, on the Objaverse benchmark (Deitke et al., 2023), Enel-7B achieves a GPT score of 50.98% for 3D object captioning, setting a new SOTA. In traditional metrics, Sentence-BERT and SimCSE reach 48.65% and 49.36%, respectively, comparable to ShapeLLM-13B. For 3D object classification, Enel-7B outperformes prior encoder-based 3D LMMs with a GPT score of 55.10%. Given the same training dataset as PointLLM, these results validate the effectiveness of our proposed LLM-embedded Semantic Encoding and Hierarchical Geometry Aggregation strategies for the encoder-free architecture. Additionally, on the 3D-VQA task of the 3D MM-Vet dataset (Qi et al., 2024), despite the lack of spatial and embodied interaction-related data in the training set, Enel achieves the GPT score of 43.1%, surpassing PointLLM-7B by 1.9%. Replacing 7B Vicuna with 13B Vicuna, Enel-13B achieves substantial performance gains across tasks. Details of the evaluation metric is provided in Appendix B.2.
@@
Response Visualization. In the Figure 6, we present a visualization of Enel’s responses for both captioning and question answering (QA) formats. We observe that in the captioning task, Enel can even accurately identify fine-grained categories such as a cricket. Moreover, in the QA task, Enel effectively handles visual challenges such as general object recognition (e.g., reasoning about numbers on dice) and spatial reasoning (e.g., accurately interpreting building structures).
@@ 5 Conclusion
Appendix A Experimental Settings
Implementation Details. We use the LLaMA model (Touvron et al., 2023) as our LLM backbone, with the 7B Vicuna-v1.1 (Chiang et al., 2023) checkpoint as the default setting. In the token embedding layer, the point cloud is first processed by a linear layer to expand its dimension from 6 to 288. The input point cloud initially consists of 8192 points, followed by three iterations of farthest point sampling (FPS), reducing the size to 512, 256, and 128, respectively. After each FPS operation, k-Nearest Neighbors (k-NN) is applied with a cluster size of 81. And geometric features are extracted using triangular encoding, followed by linear layers that progressively increase the dimension to 576, 1152, and 2304. Finally, the projection layer maps the features to the LLM dimension of 4096. In the pre-training stage, we unfreeze the first four LLM layers. Within the LLM-embedded Semantic Encoding strategy, Hybrid Semantic Loss applies masked modeling to 30% of the tokens and reconstructs the patches for the remaining 70% visible tokens. During instruction tuning, geometric aggregation is applied at the end of the 1st, 2nd, and 3rd LLM layers, reducing point tokens from 128 to 64, 32, and 16. MaxMean pooling is used to retain more information. After two LLM layers, geometric propagation is applied at the end of the 6th, 7th, and 8th layers to restore the number of point cloud to 128.
@@ Appendix A Experimental Settings
Table 6: Ablation Experiments. We begin the ablation experiments by changing the single configuration of the module from ENEL. Psi and Phi denote mask ratios of 60% and 30%, respectively. For Hybrid Semantic Loss, the subscripts patch and feat refer to the masked modeling target, with reconstruction targeting the corresponding feat and patch. $l$ represents the number of aggregation and propagation operations. $H$ refers to the LLM layers between $l$ aggregation and $l$ propagation operations. $O$ refers to the LLM layer between two individual aggregation or propagation operations.
Model Cap Cls (Avg)
@@
ENEL-7B 50.98 48.65 49.36 3.85 7.20 12.60 55.10
–Hybrid Semantic Loss 47.15 48.06 48.31 3.40 7.43 11.84 50.50
Hybrid Semantic LosspatchΦ 49.13 48.80 49.20 3.99 7.20 12.38 52.30
Hybrid Semantic LosspatchΨ 48.79 48.30 49.00 3.65 6.90 11.98 52.10
Hybrid Semantic LossfeatΨ 49.62 48.00 48.67 3.78 6.82 12.33 51.50
–gate mechanism 49.61 48.41 48.97 3.79 7.12 12.48 53.60
$l$=2, $H$=2, $O$=0 48.83 48.20 48.53 3.72 6.89 12.01 51.50
$l$=2, $H$=4, $O$=0 49.05 48.47 48.62 3.65 7.10 12.31 52.20
$l$=2, $H$=2, $O$=2 48.96 47.95 48.88 3.80 7.05 12.55 52.00
$l$=2, $H$=4, $O$=2 49.68 48.70 48.85 3.84 7.56 12.76 53.10
Appendix B More Experiments
B.1 Variants of Point Cloud Self-Supervised Learning Losses.
@@ Appendix B More Experiments
B.2 Metric Analysis
GPT-4 Evaluation is a LLM-as-a-judge framework based on custom prompts. Given a model-generated description and human reference, GPT-4 identifies key attributes from the reference, measures how many are accurately or partially matched in the model output, and returns a score from 0 to 100 with a brief explanation. It offers more comprehensive and human-aligned evaluation.
Traditional metrics like BLEU measure n-gram precision, ROUGE-L uses longest common subsequence, and METEOR combines unigram precision and recall with lemmatization and synonym matching. However, these metrics struggle with semantic similarity and tend to favor shorter outputs.
Reasons for low traditional metrics: 3D-LLM with high traditional metric scores generates captions averaging 20 words—much shorter than Enel and other methods. However, this does not indicate better output quality and performs worse in human evaluations. Traditional metrics often fail to assess the quality of detailed LLM outputs, as they favor shorter responses and struggle to capture semantic similarity. The GPT-4 score offers stronger semantic understanding, greater diversity, and better generalization.
Examples: Here is a typical example where GPT-4 gives high scores but traditional metrics give low scores. Given a point cloud of an airplane, the model outputs:
“The 3D model portrays a white cartoon airplane, styled in a simplistic and charming fashion… This model can be inferred to be used in animated children’s media or as a playful element in a game or learning application design.”
The ground truth:
“This 3D object is an airplane with distinct wings and a tail. It has a long fuselage with glass windows at the front and sides. The round-shaped wings are located in the middle.”
The model correctly identifies the object as an airplane and captures key style features like simplicity, cartoon form, and whiteness. It also reasonably infers use in children’s media, showing strong understanding. However, traditional metrics rely on n-gram overlaps. Phrases like “airplane body and wings” differ from the ground truth “fuselage with glass windows,” leading to mismatches. The output is also longer and more descriptive, while the ground truth is concise and factual, and includes extra details like “white cartoon airplane,” all contributing to low traditional scores.
B.3 More Ablation Experiments
@@
We begin the ablation experiments starting from the Enel-7B, which is the reverse order compared to the experiments in the main text, as showcased in Table 6
@@
The Effects of LLM-embedded Semantic Encoding Strategy. In the Table 6, on the basis of Enel, removing the Hybrid Semantic Loss during the pre-training stage significantly degrades performance. The GPT-4 score for the captioning task drops from 50.98% to 47.15%, and the GPT-4 score for the classification task decreases to 50.50%. This is because the proposed self-supervised learning function for point clouds effectively captures the detailed structures and high-level semantics.
@@
Based on Enel-7B, we find that setting the mask ratio in the Hybrid Semantic Loss to 30% consistently yields better results than 60%. Additionally, the configuration where the masked token part predicts features while the visible token part reconstructs patches outperforms the reverse setting—where the masked token part predicts patches and the visible token part reconstructs features. This phenomenon can be explained as follows: a mask ratio of 30% retains critical information while facilitating the model to effectively utilize the visible tokens to derive the masked parts. When the mask ratio is set too high, the model fails to learn the global context knowledge adequately. Moreover, when the masked token part is tasked with predicting features, the model focuses on learning the high-level context semantics, while the patch reconstruction aids in accurately capturing low-level details. In contrast, when the masked token part predicts patches, the model becomes excessively dependent on local features during the process of semantic reconstruction.
@@
The Effects of Hierarchical Geometry Aggregation Strategy. After removing the gating mechanism in the self-attention of the aggregation operation, the performance drops to 49.61% and 53.60% on the captioning and classification tasks, respectively. The gating mechanism helps the model to adaptively filter information, allowing it to focus on more discriminative features. Without the dynamic adjustment to focus on different parts of the input, the generated text from the LLM lacks accuracy and coherence, leading to a decrease in performance.
@@
As the number of aggregation and propagation operations decreases, overall performance declines, mainly due to insufficient layers failing to adequately model complex spatial relationships in point clouds. We observe that increasing the number of LLM layers between the final aggregation operation and the first propagation operation leads to improved performance. This suggests that fewer cascaded aggregation operations require deeper network architectures for high-level feature abstraction; otherwise, insufficient depth may lead to degraded hierarchical representations. Furthermore, the presence of LLM layers between each aggregation or propagation operation enhances performance by allowing the model to process and transform compressed information. Through self-attention mechanisms, these intermediate layers can recapture and restore details lost during the aggregation process.
Appendix C Model Output